### v1

# Hybrid Autoregressive-Diffusion Model for Real-Time Streaming Sign Language Production

Abstract
Earlier Sign Language Production (SLP) models typically relied on autoregressive methods that generate output tokens one by one, which inherently provide temporal alignment. Although techniques like Teacher Forcing can prevent model collapse during training, they still cannot solve the problem of error accumulation during inference, since ground truth is unavailable at that stage. In contrast, more recent approaches based on diffusion models leverage step-by-step denoising to enable high-quality generation. However, the iterative nature of these models and the requirement to denoise entire sequences limit their applicability in real-time tasks like SLP. To address it, we apply a hybrid approach combining autoregressive and diffusion models to SLP for the first time, leveraging the strengths of both models in sequential dependency modeling and output refinement. To capture fine-grained body movements, we design a Multi-Scale Pose Representation module that separately extracts detailed features from distinct articulators and integrates them via a Multi-Scale Fusion module. Furthermore, we introduce a Confidence-Aware Causal Attention mechanism that utilizes joint-level confidence scores to dynamically guide the pose generation process, improving accuracy and robustness. Extensive experiments on the PHOENIX14T and How2Sign datasets demonstrate the effectiveness of our method in both generation quality and real-time streaming efficiency. Project page is available at: https://hybridsign.github.io/
Introduction
Sign language serves as a vital communication medium for the deaf and hard-of-hearing communities, relying heavily on precise and expressive body and hand movements and facial expressions. The automatic generation of sign language poses significant challenges due to the complexity and subtlety of human gestures, the high-dimensional nature of pose sequences, and the need for real-time performance in practical applications. Existing approaches typically employ either autoregressive models atmc ; ptslp ; gsen , which excel at capturing temporal dependencies but may suffer from error accumulation, or diffusion models Huang2021TowardsFA ; g2pddm ; gcdm , which generate high-quality samples but are often computationally intensive and less suited for streaming scenarios.
In this work, inspired by the concept of self-forcing selfforing , we design a Hybrid Autoregressive-Diffusion Model for SLP that combines the strengths of both paradigms to enable real-time streaming generation with high fidelity and temporal coherence. Specifically, we do not adopt the teacher forcing strategy during training—instead, the model relies on its own predictions rather than ground truth data, thereby avoiding the inconsistency between training and inference.
To effectively capture the complex structure of sign language poses, we adopt a Multi-Scale Pose Representation module. This module decomposes the human pose into multiple articulators: hands, face and body, and extracts fine-grained features independently at different articulators. These features are subsequently integrated through a Multi-Scale Fusion module to form a comprehensive representation that preserves both local details and global context. This design enables our model to better handle subtle finger movements as well as facial expressions, crucial for accurate pose generation.
Another contribution is the first application of the Confidence-Aware Causal Attention mechanism to SLP. Recognizing that pose estimations often come with varying confidence levels across different joints, our attention module dynamically weighs joint contributions during generation. By incorporating joint-level confidence scores as guiding signals, the model selectively focuses on more reliable pose information, and the experiment results demonstrate that this approach leads to improved accuracy in the generated sequences. To the best of our knowledge, this is the first time that confidence information is incorporated into a causal attention mechanism for use in SLP.
Our extensive experiments on the PHOENIX14T phoenix and How2Sign how2sign sign language datasets demonstrate that the proposed model achieves the effectiveness in terms of generation quality, temporal consistency, and computational efficiency. Moreover, the streaming capability of our hybrid model makes it highly practical for real-world sign language communication systems, including live interpretation and interactive applications.
Our main contributions are summarized as follows:
  • We are the first to apply a hybrid autoregressive-diffusion framework to SLP, effectively balancing temporal modeling and high-quality refinement for real-time generation.
  • We design a Multi-Scale Pose Representation module and a Multi-Scale Fusion module, which effectively capture fine-grained articulator features and their interactions.
  • We are the first to introduce the Confidence-Aware Causal Attention mechanism to SLP, leveraging joint-level confidence scores to enhance generation accuracy and robustness.
The rest of this paper is organized as follows: Section 2 reviews related work, Section 3 outlines the preliminary, Section 4 details the proposed model architecture and components, Section 5 presents experimental results, and Section 6 concludes with future directions.
Related Work
Sign language is the primary means of communication for deaf and non-deaf people. However, most hearing individuals have not learned sign language, resulting in significant communication barriers between the two groups. Early research efforts mainly focused on Sign Language Recognition (SLR) slr1 ; slr2 ; slr3 ; slr4 ; slr5 and Sign Language Translation (SLT) slt1 ; slt2 ; slt3 . However, with the advancement of technology and the growing demand for more efficient communication, increasing attention has been directed toward Sign Language Production (SLP).
Sign Language Production (SLP)
Initial studies in SLP predominantly relied on rule-based animation techniques to translate textual inputs into synthetic avatar-based sign language animations slp1 ; slp2 ; slp3 . These methods typically employed predefined templates and handcrafted lookup rules to map sentences or glosses to gestures. While these systems offered interpretable sign animations, they suffered from high data collection costs, limited scalability, and poor generalization to unseen sentences or grammatical structures.
Autoregressive and Diffusion models in SLP
With the advent of deep learning, more flexible and data-driven SLP models were introduced. Text2Sign text2sign introduces a multi-stage pipeline for sign language production (SLP), which decomposes the overall process into three sequential subtasks: text-to-gloss [^1] Gloss is a written representation of signs using capitalized words to show the meaning and order of the signs. translation, gloss-to-pose generation, and pose-to-video synthesis. While effective, such pipeline designs introduced potential efficiency issues and error propagation between stages. To address this, ptslp presented the first end-to-end autoregressive model that directly generates sign pose sequences from glosses. Further improvements followed, including a multi-channel adversarial model atmc and a Mixture Density Transformer 3dmc to better model the multi-modal nature of sign gestures.
Recently, diffusion models have shown strong potential in SLP due to their capacity for modeling complex output distributions. Methods such as SignDiff signdiff , G2P-DDM g2pddm , and GCDM gcdm leverage denoising diffusion probabilistic models (DDPMs ddpm ) to generate sign pose sequences by gradually refining Gaussian noise into meaningful joint coordinates. These approaches typically treat pose generation as a coordinate regression problem, guided by gloss or semantic input. Subsequently, Sign-IDD signidd integrates an attribute-controllable diffusion module that leverages skeletal direction and length attributes to constrain joint associations, thereby enabling more precise and controllable pose generation.
Although both autoregressive and diffusion-based models have achieved significant progress, real-time sign language production remains highly challenging. Autoregressive models naturally excel at temporally modeling and can output each frame immediately without waiting for the entire sequence to be generated, making them highly suitable for real-time scenarios. However, even with strategies like Teacher Forcing to mitigate training collapse, these models still suffer from exposure bias during inference due to their reliance on predicted frames rather than ground truth. On the other hand, diffusion models produce high-quality generation but rely on an iterative denoising process that introduces substantial latency, limiting their applicability in real-time settings. To overcome these limitations, we for the first time apply a hybrid model to SLP that leverages the advantages of both approaches. Our method is able to generate the first frame of the video in 1.22 seconds and then stream subsequent high-quality frames continuously. Compared to the methods we are aware of, it takes significantly less time and effectively addresses the challenges of real-time sign language production.
Preliminary
In this section, we introduce the fundamental concepts and notations used in our framework, including autoregressive modeling, diffusion-based generation, human pose representation, and attention mechanisms.
Autoregressive Modeling
Autoregressive (AR) models generate a sequence by modeling each element conditioned on its previous elements. For a sequence of human poses $P=\{P_{1},P_{2},\ldots,P_{T}\}$, where $P_{t}\in R^{J\times D}$ represents the positions of $J$ joints at time $t$ in $D$-dimensional space, the autoregressive factorization is:
$p(P)=\prod_{t=1}^{T}p(P_{t}\mid P_{<t})$ (1)
This formulation allows for temporal modeling but is prone to error accumulation due to its greedy generation process.
Diffusion Models
Score-based diffusion models.
Diffusion models learn to generate data by reversing a gradual noising process. Tranditional diffusion models sde ; ddim ; iddpm follow the denoising diffusion probabilistic model (DDPM) framework ddpm , where a data sample $x_{0}$ is progressively perturbed over $T$ steps using a predefined noise schedule $\{\beta_{t}\}_{t=1}^{T}$. The forward process is defined as:
$q(x_{t}\mid x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (2)
This results in:
$q(x_{t}\mid x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I)$ (3)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$.
The denoising model is trained to predict the added noise $\boldsymbol{\epsilon}$:
$\mathcal{L}_{\text{Diff}}=\mathbb{E}_{x_{0},t,\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(x_{t},t)\right\|^{2}\right]$ (4)
However, traditional diffusion models have slow sampling speeds, which significantly hinder real-time generation.
Flow-based Diffusion Models.
In contrast to DDPMs, flow-based diffusion models flowmatching1 ; flowmatching2 ; rectifiedflow learn an explicit transformation between noise and data using invertible mappings. These models parameterize the denoising process as an ODE (or deterministic flow) rather than a stochastic process. Specifically, they model a continuous-time trajectory between the data distribution and a base noise distribution using a neural network as the vector field.
The process is defined through the following differential equation:
$\frac{dx(t)}{dt}=v_{\theta}(x(t),t),\quad x(0)\sim\mathcal{N}(0,I)$ (5)
Here, $v_{\theta}$ is a neural network trained to match the score or displacement between perturbed data and clean data. This flow transports samples from the base distribution (e.g., Gaussian) at $t=0$ to the data distribution at $t=1$.
A popular training objective is the flow matching loss:
$\mathcal{L}_{\text{Flow}}=\mathbb{E}_{x_{0},t,z}\left[\left|v_{\theta}(x_{t},t)-\frac{x_{0}-x_{t}}{t}\right|^{2}\right]$ (6)
where $x_{t}=(1-t)z+tx_{0}$ is a linear interpolation between a noise sample $z\sim\mathcal{N}(0,I)$ and a data sample $x_{0}$.
This objective encourages the model to learn a displacement field that transports noise into data along straight paths, offering a deterministic alternative to stochastic diffusion, thereby accelerating training efficiency and sampling speed.
Human Pose Representation
A human pose at time $t$ is represented as a set of $J$ joints:
$P_{t}=\{p_{t,1},p_{t,2},\ldots,p_{t,J}\},\quad p_{t,j}\in\mathbb{R}^{D}$ (7)
Each joint is also associated with a confidence score $c_{t,j}\in[0,1]$, denoted as:
$C_{t}=\{c_{t,1},c_{t,2},\ldots,c_{t,J}\}$ (8)
These confidence scores are used later in our attention mechanism to modulate the influence of each joint during pose generation (Section 4.3).
Attention Mechanism
We build upon the standard scaled dot-product attention mechanism. Given query $Q\in\mathbb{R}^{n\times d}$, key $K\in\mathbb{R}^{m\times d}$, and value $V\in\mathbb{R}^{m\times d}$ matrices, the attention is computed as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V$ (9)
To preserve temporal causality in streaming generation, we apply a causal mask to prevent attention to future time steps.
Later, in our proposed method, we extend this formulation to include confidence-awareness, biasing the attention distribution based on joint reliability (Section 4.3).
Figure 1: Overview of Hybrid Autoregressive-Diffusion Model.
Methodology
Our proposed framework for real-time streaming sign language production integrates three key components:
  • a Hybrid Autoregressive-Diffusion framework,
  • a Multi-Scale Pose Representation module and Fusion module, and
  • a Confidence-Aware Causal Attention mechanism.
These components work together to ensure temporally coherent and high-quality sign language pose generation.
Figure 2: Expert module for facial features in the Multi-Scale Pose Representation Module.
Overview
As shown in Figure 1, given a natural language sentence (e.g., in German or English), the model generates a temporally consistent sequence of human poses corresponding to sign language gestures in a streaming, frame-by-frame manner. First, the Multi-Scale Pose Representation module uses dedicated Expert modules to generate three articulators: face, body, and hands, through a diffusion-based process. To support real-time generation of full motion frames, these articulators are fused into a complete pose frame by the Multi-Scale Fusion module. The fused pose is then decomposed back into the three articulators. Each part serves as a conditional input for generating the corresponding part in the next frame, following an autoregressive process. The entire generation pipeline is guided by joint confidence scores to enhance both the accuracy and realism of the output.
Hybrid Autoregressive-Diffusion Model
We apply, for the first time, a hybrid generative model that integrates autoregressive and diffusion paradigms to SLP, aiming to harness the advantages of both.
Specifically, we introduce causal attention into the denoiser of the diffusion model by applying an attention mask $M\in\mathbb{R}^{n\times n}$ to its self-attention mechanism:
$M_{ij}=\begin{cases}0&\text{if }j\leq i\\ -\infty&\text{if }j>i\end{cases}$ (10)
so that each position $i$ can only attend to positions $j\leq i$ (i.e., current and previous tokens).
Figure 3: Confidence-Aware Causal Attention Mechanism.
To address the distribution mismatch between training and inference commonly observed in autoregressive models, we draw on the training strategy proposed in Self-Forcing selfforing : instead of relying on ground truth during training, our model autoregressively generates each frame by first generating its articulators: face, hands and body, based on previously generated outputs. Specifically, at each time step $t$, the model produces:
$\tilde{C}_{t}$ $=f_{\text{AR}}(\tilde{C}_{<t}),$ (11)
$\text{with}\quad\tilde{C}_{<t}$ =$=\{\tilde{C}_{1},\dots,\tilde{C}_{t-1}\},\quad\tilde{C}_{t}=\{\tilde{F}_{t},\tilde{H}_{t},\tilde{B}_{t}\}$
where $\tilde{F}_{t}$, $\tilde{H}_{t}$, and $\tilde{B}_{t}$ represent the generated face, hands, and body components at time step $t$, respectively. These components are then fused to obtain the final full-frame pose representation. This encourages the model to develop robustness against its own generation errors and better aligns the training process with inference. Moreover, we adopt holistic loss—soft-DTW softdtw —which directly evaluates the quality of the entire generated sequence, thereby enabling more consistent supervision and further mitigating error accumulation. For further details, please refer to Section 4.4.
To improve training efficiency and sampling speed, we adopt a flow-based diffusion model flowmatching1 ; flowmatching2 instead of the traditional DDPM-based approach ddpm ; ddim . Flow matching enables learning a continuous transformation path that maps the initial distribution directly to the target distribution, allowing sampling to be completed in fewer steps and enhancing both training and sampling efficiency.
Multi-Scale Pose Representation Module
Sign language involves coordinated motion across face, body, and hands. To capture this, we design a Multi-Scale Pose Representation module (Figure 2) that partitions keypoints into three anatomical groups and processes each with a dedicated expert module.
Given a sequence of $T$ frames with $N$ keypoints per frame, each keypoint $p_{t}^{(i)}=[x_{t}^{(i)},y_{t}^{(i)},c_{t}^{(i)}]$ is embedded as:
$s_{t}^{(i)}=\mathrm{MLP}(p_{t}^{(i)})+PE(t),$ (12)
We divide the keypoints into groups $\mathcal{G}={G_{\text{face}},G_{\text{body}},G_{\text{hands}}}$ and extract group-specific features:
$S_{t}^{(k)}={s_{t}^{(i)}\mid i\in G_{k}},\quad k=1,2,3$ (13)
Each group sequence is processed by a scale-specific expert $\mathcal{E}^{(k)}$ to capture intra-scale dependencies:
$H^{(k)}=\mathcal{E}^{(k)}([S_{1}^{(k)},\dots,S_{T}^{(k)}])$ (14)
We average-pool over joints and apply attention-based fusion across scales:
$H_{t}=\sum_{k=1}^{3}\alpha_{t}^{(k)}\cdot H_{t}^{(k)},\quad\alpha_{t}^{(k)}=\frac{\exp(w^{\top}\tanh(w_{f}H_{t}^{(k)}))}{\sum_{j}\exp(w^{\top}\tanh(w_{f}H_{t}^{(j)}))}$ (15)
The fused representation $H_{t}$ serves as input for subsequent generation modules.
Confidence-Aware Causal Attention Mechanism
To improve robustness in autoregressive pose generation, we propose a Confidence-Aware Causal Attention module (Figure 3), which integrates keypoint confidence scores into the attention computation to downweight unreliable inputs.
Given a sentence $S$, the model generates a pose sequence $x={x^{(1)},\dots,x^{(T)}}$, where each $x^{(t)}\in\mathbb{R}^{J\times d}$ represents $J$ keypoints. Each keypoint has an associated confidence score $c_{i}^{(t)}\in[0,1]$.
Table 1: Performance comparison on PHOENIX14T
Methods DEV TEST
[…]
G2P-DDM g2pddm - - - - - - 16.11 7.50 - 77.26 - -
GCDM gcdm 22.88 7.64 23.35 82.81 11.18 - 22.03 7.91 23.20 81.94 11.10 -
Sign-IDD signidd 25.40 8.93 27.60 77.72 5.09 39.11 24.80 9.08 26.58 76.66 6.20 47.19
HybridSign (Ours) 26.98 9.26 28.07 75.81 4.07 38.28 25.77 10.03 27.97 75.02 4.96 45.50
Ground Truth 29.77 12.13 29.60 74.17 0.00 0.00 29.76 11.93 28.98 71.94 0.00 0.00
Table 2: Performance comparison on How2Sign
Methods DEV TEST
[…]
Sign-IDD signidd 29.12 6.27 17.01 89.95 4.76 35.03 28.90 6.06 16.21 89.98 4.86 39.02
HybridSign (Ours) 30.71 6.92 18.96 87.11 3.76 34.19 30.12 6.48 18.02 88.30 3.89 37.10
Ground Truth 35.20 8.89 22.45 83.79 0.00 0.00 34.01 8.03 21.87 81.94 0.00 0.00
Confidence-Weighted Attention.
In causal attention, timestep $t$ only attends to ${1,\dots,t}$:
$\alpha_{t,s}=\frac{\exp\left(\frac{Q_{t}K_{s}^{\top}}{\sqrt{d_{k}}}\right)}{\sum_{j=1}^{t}\exp\left(\frac{Q_{t}K_{j}^{\top}}{\sqrt{d_{k}}}\right)}$ (16)
We introduce a confidence bias to this formulation:
$\alpha_{t,s}=\frac{\exp\left(\frac{Q_{t}K_{s}^{\top}}{\sqrt{d_{k}}}+\beta\cdot\bar{c}^{(s)}\right)}{\sum_{j=1}^{t}\exp\left(\frac{Q_{t}K_{j}^{\top}}{\sqrt{d_{k}}}+\beta\cdot\bar{c}^{(j)}\right)}$ (17)
where $\bar{c}^{(s)}=\frac{1}{J}\sum_{i=1}^{J}c_{i}^{(s)}$ is the average keypoint confidence, and $\beta$ is a learnable scalar controlling the strength of the bias.
This allows the model to attend more to reliable frames during decoding.
Training Objectives
The model is trained using three loss terms:
Joint Loss.
To ensure accurate joint positions:
$\mathcal{L}_{joint}=\frac{1}{J}\sum_{j=1}^{J}|p_{j}-\hat{p}_{j}|$ (18)
where $p_{j}$ and $\hat{p}_{j}$ are ground truth and predicted joint positions.
Bone Loss.
To preserve bone orientations:
$\mathcal{L}_{bone}=\frac{1}{B}\sum_{b=1}^{B}(q_{b}-\hat{q}_{b})^{2}$ (19)
where $q_{b}$ and $\hat{q}_{b}$ denote ground truth and predicted bone orientations.
Soft-DTW Loss.
To align pose sequences temporally, we use Soft-DTW softdtw :
$\mathcal{L}_{\text{soft-dtw}}=\min\nolimits_{A}^{\gamma}\sum_{(i,j)\in A}\|x_{i}-y_{j}\|^{2}$
with the smooth minimum operator:
$\min\nolimits^{\gamma}\{a_{1},\dots,a_{k}\}=-\gamma\log\left(\sum_{i=1}^{k}e^{-a_{i}/\gamma}\right)$
Total Loss.
Dynamic weighting based on inverse EMA dwa balances all terms:
$\lambda_{i}^{(t)}=\frac{1}{\bar{\mathcal{L}}_{i}^{(t)}+\epsilon}\bigg{/}\sum_{j=1}^{N}\frac{1}{\bar{\mathcal{L}}_{j}^{(t)}+\epsilon}$ (20)
The final total loss is:
$\mathcal{L}_{\text{total}}^{(t)}=\lambda_{1}^{(t)}\mathcal{L}_{joint}^{(t)}+\lambda_{2}^{(t)}\mathcal{L}_{bone}^{(t)}+\lambda_{3}^{(t)}\mathcal{L}_{soft\text{-}dtw}^{(t)}$ (21)
This scheme balances the losses to stabilize training.
Experiment
Figure 4: Visualization examples of generated poses on How2Sign. We compare HybridSign with the ground-truth poses and the original video frames for clear evaluation.
Experimental Setup
Datasets.
The benchmark datasets we tested in our experiments are as follows:
  • The RWTH-PHOENIX-Weather 2014T dataset phoenix is a German Sign Language (DGS) corpus that consists of 8,257 sentence-level samples, encompassing 2,887 unique German words. Specifically, it includes 7,096 training samples (featuring 1,066 distinct sign glosses in the gloss annotations and 2,887 words in the corresponding German spoken language translations), 519 validation samples, and 642 test samples.
  • The How2Sign dataset how2sign is a large-scale American Sign Language (ASL) corpus designed for sign language understanding and generation tasks. It contains over 80 hours of multimodal data, including RGB video and 2D keypoints. The dataset training dataset consists of 31,048 video segments, with 1,739 segments in the development set and 2,343 segments in the test set.
Evaluation Metrics.
Following the existing works g2pddm ; gcdm ; signidd ; signd2c , we also employed four types of metrics to evaluate the model’s performance, including BLEU bleu , ROUGE rouge , WER, DTW and FID. However, since SLP cannot directly compute these metrics, we employed a pre-trained SLT model slt for back-translation. Specifically, the generated sign language videos are translated back into natural language, and the original text input is compared with the translated output to calculate these metrics.
Comparisons with Baseline Models
We conducted both qualitative and quantitative analyses to evaluate the model’s performance.
Table 3: Latency and throughput comparison between the baseline model and this model with different numbers of expert modules for generating a 5-second, 60-frame sign language pose sequence. The total time includes processing by the text encoder and the generative model.
Method Latency (s)$\downarrow$ Throughput (FPS)$\uparrow$
Sign-IDD signidd 40.31 1.49
HybridSign with 4 experts 2.02 5.66
HybridSign with 3 experts 1.22 10.17
  • Note: ”4 experts” refer to face, body, left hand, and right hand. ”3 experts” refer to face, body, and hands combined.
Quantitative Analysis.
Tables 1 and 2 present a performance comparison between the proposed HybridSign and various baseline models on the PHOENIX14Tphoenix and How2Signhow2sign datasets, respectively. As shown, HybridSign outperforms all baseline models across all evaluation metrics.
Table 3 compares latency and throughput with the baseline model. Here, latency refers to the time required to generate the first frame of the 60-frame sequence, rather than the total time to produce the entire video. This reflects the model’s initial response speed, which is important for real-time applications. HybridSign achieves significantly lower latency and higher throughput than the baseline, indicating its efficiency and practicality in real-time sign language production.
Performance differences across the two datasets are relatively small, likely due to the use of 2D pose sequences. While this may slightly reduce generation quality, it lowers computational complexity and improves training and inference efficiency. To compensate, we incorporate the Multi-Scale Pose Representation Module and the Confidence-Attention Causal Mechanism, which enhance overall performance. Notably, the DTW score is reduced by approximately 80% thanks to the Soft-DTW loss softdtw , which improves temporal alignment.
Qualitative Analysis.
Figure 4 illustrates examples of sign poses generated by HybridSign alongside the corresponding ground truth. While minor differences are observable, the generated poses remain highly similar overall and do not hinder interpretability. This further supports the effectiveness and practicality of our approach.
From the generated results, there are still some subtle differences compared to the ground truth, such as variations in bone lengths or joint bending angles. However, based on the results shown in the table 1 and 2, after performing back translation on the sign pose sequences, the B1 and B4 scores on the two datasets are 26.98, 9.26 and 30.71, 6.92 respectively. Therefore, we believe these slight differences are acceptable.
Table 4: Ablation study results of modules on the How2Sign dataset.
Modules DEV TEST
[…]
RNN 23.47 5.02 6.89 22.92 4.72 7.84
Causal Attention 26.46 5.71 5.10 25.08 5.83 5.50
Confidence-Aware Causal Attention 30.71 6.92 3.76 30.12 6.48 3.89
Table 5: Ablation study results of expert numbers on the How2Sign dataset.
Export Numbers DEV TEST
[…]
1 (whole pose) 22.73 5.31 6.85 22.33 5.14 7.02
4 (face + body + lh + rh) 30.83 6.98 3.72 30.31 6.56 3.77
3 (face + body + hands) 30.71 6.92 3.76 30.12 6.48 3.89
Ablation Study
We also conducted two ablation studies to evaluate the effects of using different attention mechanisms and varying numbers of expert modules.
Table 4 shows the performance differences among various autoregressive implementations. The model using Confidence-Aware Causal Attention outperforms those using RNN rnn or standard Causal Attention attention . This is because the attention mechanism takes confidence scores into account, assigning different attention weights to keypoints based on their reliability.
Table 5 presents the results of varying the number of expert modules. When only a single expert module is used—treating the entire pose as a whole—the generation quality drops significantly due to the inability to capture local details. Using three or four expert modules yields similar performance, but adding more experts increases the number of parameters, leading to higher computational cost 3. Therefore, we choose to use three expert modules as a balanced trade-off between performance and efficiency.
Conclusions
This paper is the first to apply a hybrid framework that combines autoregressive and diffusion models to sign language production, enabling real-time, streaming generation of sign language pose frames. The framework leverages the inherent sequential modeling capability and frame-by-frame output of autoregressive models, along with the high-quality generation strengths of diffusion models. Additionally, it incorporates multi-scale pose representation modeling and a confidence-aware causal attention mechanism, which further improve the naturalness and expressive accuracy of the generated sign language. Extensive experiments on several benchmark datasets demonstrate the effectiveness of the proposed method, especially in real-time application scenarios.