### v3

# Hybrid Autoregressive-Diffusion Model for Real-Time Sign Language Production

Abstract
Earlier Sign Language Production (SLP) models typically relied on autoregressive methods that generate output tokens one by one, which inherently provide temporal alignment. Although techniques like Teacher Forcing can prevent model collapse during training, they still cannot solve the problem of error accumulation during inference, since ground truth is unavailable at that stage. In contrast, more recent approaches based on diffusion models leverage step-by-step denoising to enable high-quality generation. However, the iterative nature of these models and the requirement to denoise entire sequences limit their applicability in real-time tasks like SLP. To address it, we explore a hybrid approach that combines autoregressive and diffusion models for SLP, leveraging the strengths of both models in sequential dependency modeling and output refinement. To capture fine-grained body movements, we design a Multi-Scale Pose Representation module that separately extracts detailed features from distinct articulators and integrates them via a Multi-Scale Fusion module. Furthermore, we introduce a Confidence-Aware Causal Attention mechanism that utilizes joint-level confidence scores to dynamically guide the pose generation process, improving accuracy and robustness. Extensive experiments on the PHOENIX14T and How2Sign datasets demonstrate the effectiveness of our method in both generation quality and real-time efficiency.
Introduction
Sign language serves as a vital communication medium for the deaf and hard-of-hearing communities, relying heavily on precise and expressive body and hand movements and facial expressions. The automatic generation of sign language poses significant challenges due to the complexity and subtlety of human gestures, the high-dimensional nature of pose sequences, and the need for real-time performance in practical applications. Existing approaches typically employ either autoregressive models atmc ; ptslp ; gsen , which excel at capturing temporal dependencies but may suffer from error accumulation, or diffusion models Huang2021TowardsFA ; g2pddm ; gcdm , which generate high-quality samples but are often computationally intensive and less suited for real-time scenarios.
In this work, inspired by the concept of self-forcing selfforing , we design a Hybrid Autoregressive-Diffusion Model for SLP that combines the strengths of both paradigms to enable real-time generation with high fidelity and temporal coherence. Specifically, we do not adopt the teacher forcing strategy during training—instead, the model relies on its own predictions rather than ground truth data, thereby avoiding the inconsistency between training and inference.
To effectively capture the complex structure of sign language poses, we adopt a Multi-Scale Pose Representation module. This module decomposes the human pose into multiple articulators: hands, face and body, and extracts fine-grained features independently at different articulators. These features are subsequently integrated through a Multi-Scale Fusion module to form a comprehensive representation that preserves both local details and global context. This design enables our model to better handle subtle finger movements as well as facial expressions, crucial for accurate pose generation.
Another contribution is the first application of the Confidence-Aware Causal Attention mechanism to SLP. Recognizing that pose estimations often come with varying confidence levels across different joints, our attention module dynamically weighs joint contributions during generation. By incorporating joint-level confidence scores as guiding signals, the model selectively focuses on more reliable pose information, and the experiment results demonstrate that this approach leads to improved accuracy in the generated sequences. To the best of our knowledge, this is the first time that confidence information is incorporated into a causal attention mechanism for use in SLP.
Our extensive experiments on the PHOENIX14T phoenix and How2Sign how2sign datasets demonstrate that the proposed model delivers high generation quality, strong temporal consistency, and notable computational efficiency, making it highly practical for real-world sign language communication systems, including live interpretation and interactive applications.
Our main contributions are summarized as follows:
  • We are the first to apply a hybrid autoregressive-diffusion framework to SLP, effectively balancing temporal modeling and high-quality refinement for real-time generation.
  • We design a Multi-Scale Pose Representation module and a Multi-Scale Fusion module, which effectively capture fine-grained articulator features and their interactions.
  • We are the first to introduce the Confidence-Aware Causal Attention mechanism to SLP, leveraging joint-level confidence scores to enhance generation accuracy and robustness.
The rest of this paper is organized as follows: Section 2 reviews related work, Section 3 outlines the preliminary, Section 4 details the proposed model architecture and components, Section 5 presents experimental results, and Section 6 concludes with future directions.
Related Work
Sign language is the primary means of communication for deaf and non-deaf people. However, most hearing individuals have not learned sign language, resulting in significant communication barriers between the two groups. Early research efforts mainly focused on Sign Language Recognition (SLR) slr1 ; slr2 ; slr3 ; slr4 ; slr5 and Sign Language Translation (SLT) slt1 ; slt2 ; slt3 . However, with the advancement of technology and the growing demand for more efficient communication, increasing attention has been directed toward Sign Language Production (SLP).
Sign Language Production (SLP)
Initial studies in SLP predominantly relied on rule-based animation techniques to translate textual inputs into synthetic avatar-based sign language animations slp1 ; slp2 ; slp3 . These methods typically employed predefined templates and handcrafted lookup rules to map sentences or glosses to gestures. While these systems offered interpretable sign animations, they suffered from high data collection costs, limited scalability, and poor generalization to unseen sentences or grammatical structures.
Autoregressive and Diffusion models in SLP
With the advent of deep learning, more flexible and data-driven SLP models were introduced. Text2Sign text2sign introduces a multi-stage pipeline for sign language production (SLP), which decomposes the overall process into three sequential subtasks: text-to-gloss [^1] Gloss is a written representation of signs using capitalized words to show the meaning and order of the signs. translation, gloss-to-pose generation, and pose-to-video synthesis. While effective, such pipeline designs introduced potential efficiency issues and error propagation between stages. To address this, ptslp presented the first end-to-end autoregressive model that directly generates sign pose sequences from glosses. Further improvements followed, including a multi-channel adversarial model atmc and a Mixture Density Transformer 3dmc to better model the multi-modal nature of sign gestures.
Recently, diffusion models have shown strong potential in SLP due to their capacity for modeling complex output distributions. Methods such as SignDiff signdiff , G2P-DDM g2pddm , and GCDM gcdm leverage denoising diffusion probabilistic models (DDPMs ddpm ) to generate sign pose sequences by gradually refining Gaussian noise into meaningful joint coordinates. These approaches typically treat pose generation as a coordinate regression problem, guided by gloss or semantic input. Subsequently, Sign-IDD signidd integrates an attribute-controllable diffusion module that leverages skeletal direction and length attributes to constrain joint associations, thereby enabling more precise and controllable pose generation.
Although both autoregressive and diffusion-based models have achieved significant progress, real-time sign language production remains highly challenging. Autoregressive models naturally excel at temporally modeling and can output each frame immediately without waiting for the entire sequence to be generated, making them highly suitable for real-time scenarios. However, even with strategies like Teacher Forcing to mitigate training collapse, these models still suffer from exposure bias during inference due to their reliance on predicted frames rather than ground truth. On the other hand, diffusion models produce high-quality generation but rely on an iterative denoising process that introduces substantial latency, limiting their applicability in real-time settings. To overcome these limitations, we for the first time apply a hybrid model to SLP that leverages the advantages of both approaches. Our method is able to generate the first frame of the video in 5.9 seconds and then subsequent high-quality frames continuously. Compared to the methods we are aware of, it takes significantly less time and effectively addresses the challenges of real-time sign language production.
Preliminary
In this section, we introduce the fundamental concepts and notations used in our framework, including autoregressive modeling, diffusion-based generation, human pose representation, and attention mechanisms.
Autoregressive Modeling
Autoregressive models generate a sequence by modeling each element conditioned on its previous elements. For a sequence of human poses $P=\{P_{1},P_{2},\ldots,P_{T}\}$, where $P_{t}\in R^{J\times D}$ represents the positions of $J$ joints at time $t$ in $D$-dimensional space, the autoregressive factorization is:
$p(P)=\prod_{t=1}^{T}p(P_{t}\mid P_{<t})$ (1)
This formulation allows for temporal modeling but is prone to error accumulation due to its greedy generation process.
Diffusion Models
Score-based diffusion models.
Diffusion models learn to generate data by reversing a gradual noising process. Tranditional diffusion models sde ; ddim ; iddpm follow the denoising diffusion probabilistic model (DDPM) framework ddpm , where a data sample $x_{0}$ is progressively perturbed over $T$ steps using a predefined noise schedule $\{\beta_{t}\}_{t=1}^{T}$. The forward process is defined as:
$q(x_{t}\mid x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (2)
This results in:
$q(x_{t}\mid x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I)$ (3)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$.
The denoising model is trained to predict the added noise $\boldsymbol{\epsilon}$:
$\mathcal{L}_{\text{Diff}}=\mathbb{E}_{x_{0},t,\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(x_{t},t)\right\|^{2}\right]$ (4)
However, traditional diffusion models have slow sampling speeds, which significantly hinder real-time generation.
Flow-based Diffusion Models.
In contrast to DDPMs, flow-based diffusion models flowmatching1 ; flowmatching2 ; rectifiedflow learn an explicit transformation between noise and data using invertible mappings. These models parameterize the denoising process as an ODE (or deterministic flow) rather than a stochastic process. Specifically, they model a continuous-time trajectory between the data distribution and a base noise distribution using a neural network as the vector field.
The process is defined through the following differential equation:
$\frac{dx(t)}{dt}=v_{\theta}(x(t),t),\quad x(0)\sim\mathcal{N}(0,I)$ (5)
Here, $v_{\theta}$ is a neural network trained to match the score or displacement between perturbed data and clean data. This flow transports samples from the base distribution (e.g., Gaussian) at $t=0$ to the data distribution at $t=1$.
A popular training objective is the flow matching loss:
$\mathcal{L}_{\text{Flow}}=\mathbb{E}_{x_{0},t,z}\left[\left|v_{\theta}(x_{t},t)-\frac{x_{0}-x_{t}}{t}\right|^{2}\right]$ (6)
where $x_{t}=(1-t)z+tx_{0}$ is a linear interpolation between a noise sample $z\sim\mathcal{N}(0,I)$ and a data sample $x_{0}$.
This objective encourages the model to learn a displacement field that transports noise into data along straight paths, offering a deterministic alternative to stochastic diffusion, thereby accelerating training efficiency and sampling speed.
Human Pose Representation
A human pose at time $t$ is represented as a set of $J$ joints:
$P_{t}=\{p_{t,1},p_{t,2},\ldots,p_{t,J}\},\quad p_{t,j}\in\mathbb{R}^{D}$ (7)
Each joint is also associated with a confidence score $c_{t,j}\in[0,1]$, denoted as:
$C_{t}=\{c_{t,1},c_{t,2},\ldots,c_{t,J}\}$ (8)
These confidence scores are used later in our attention mechanism to modulate the influence of each joint during pose generation (Section 4.3).
Attention Mechanism
We build upon the standard scaled dot-product attention mechanism. Given query $Q\in\mathbb{R}^{n\times d}$, key $K\in\mathbb{R}^{m\times d}$, and value $V\in\mathbb{R}^{m\times d}$ matrices, the attention is computed as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V$ (9)
To preserve temporal causality in generation, we apply a causal mask to prevent attention to future time steps.
Later, in our proposed method, we extend this formulation to include confidence-awareness, biasing the attention distribution based on joint reliability (Section 4.3).
Figure 1: Overview of Hybrid Autoregressive-Diffusion Model.
Methodology
Our proposed framework for real-time sign language production integrates three key components:
  • a Hybrid Autoregressive-Diffusion framework,
  • a Multi-Scale Pose Representation module and Fusion module, and
  • a Confidence-Aware Causal Attention mechanism.
These components work together to ensure temporally coherent and high-quality sign language pose generation.
Figure 2: Expert module for facial features in the Multi-Scale Pose Representation Module.
Overview
As shown in Figure 1, given a natural language sentence (e.g., in German or English), the model generates a temporally consistent sequence of human poses corresponding to sign language gestures in a frame-by-frame manner. First, the Multi-Scale Pose Representation module uses dedicated Expert modules to generate three articulators: face, body, and hands, through a diffusion-based process. To support real-time generation of full motion frames, these articulators are fused into a complete pose frame by the Multi-Scale Fusion module. The fused pose is then decomposed back into the three articulators. Each part serves as a conditional input for generating the corresponding part in the next frame, following an autoregressive process. The entire generation pipeline is guided by joint confidence scores to enhance both the accuracy and realism of the output.
Hybrid Autoregressive-Diffusion Model
We apply, for the first time, a hybrid generative model that integrates autoregressive and diffusion paradigms to SLP, aiming to harness the advantages of both.
Specifically, we introduce causal attention into the denoiser of the diffusion model by applying an attention mask $M\in\mathbb{R}^{n\times n}$ to its self-attention mechanism:
$M_{ij}=\begin{cases}0&\text{if }j\leq i\\ -\infty&\text{if }j>i\end{cases}$ (10)
so that each position $i$ can only attend to positions $j\leq i$ (i.e., current and previous tokens).
Figure 3: Confidence-Aware Causal Attention Mechanism.
To address the distribution mismatch between training and inference commonly observed in autoregressive models, we draw on the training strategy proposed in Self-Forcing selfforing : instead of relying on ground truth during training, our model autoregressively generates each frame by first generating its articulators: face, hands and body, based on previously generated outputs. Specifically, at each time step $t$, the model produces:
$\tilde{C}_{t}$ $=f_{\text{AR}}(\tilde{C}_{<t}),$ (11)
$\text{with}\quad\tilde{C}_{<t}$ $=\{\tilde{C}_{1},\dots,\tilde{C}_{t-1}\},\quad\tilde{C}_{t}=\{\tilde{F}_{t},\tilde{H}_{t},\tilde{B}_{t}\}$
where $\tilde{F}_{t}$, $\tilde{H}_{t}$, and $\tilde{B}_{t}$ represent the generated face, hands, and body components at time step $t$, respectively. These components are then fused to obtain the final full-frame pose representation. This encourages the model to develop robustness against its own generation errors and better aligns the training process with inference. Moreover, we adopt holistic loss—soft-DTW softdtw —which directly evaluates the quality of the entire generated sequence, thereby enabling more consistent supervision and further mitigating error accumulation. For further details, please refer to Section 4.4.
To improve training efficiency and sampling speed, we adopt a flow-based diffusion model flowmatching1 ; flowmatching2 instead of the traditional DDPM-based approach ddpm ; ddim . Flow matching enables learning a continuous transformation path that maps the initial distribution directly to the target distribution, allowing sampling to be completed in fewer steps and enhancing both training and sampling efficiency.
Multi-Scale Pose Representation Module
Sign language involves coordinated motion across face, body, and hands. To capture this, we design a Multi-Scale Pose Representation module (Figure 2) that partitions keypoints into three anatomical groups and processes each with a dedicated expert module.
Given a sequence of $T$ frames with $N$ keypoints per frame, each keypoint $p_{t}^{(i)}=[x_{t}^{(i)},y_{t}^{(i)},c_{t}^{(i)}]$ is embedded as:
$s_{t}^{(i)}=\mathrm{MLP}(p_{t}^{(i)})+PE(t),$ (12)
We divide the keypoints into groups $\mathcal{G}={G_{\text{face}},G_{\text{body}},G_{\text{hands}}}$ and extract group-specific features:
$S_{t}^{(k)}={s_{t}^{(i)}\mid i\in G_{k}},\quad k=1,2,3$ (13)
Each group sequence is processed by a scale-specific expert $\mathcal{E}^{(k)}$ to capture intra-scale dependencies:
$H^{(k)}=\mathcal{E}^{(k)}([S_{1}^{(k)},\dots,S_{T}^{(k)}])$ (14)
We average-pool over joints and apply attention-based fusion across scales:
$H_{t}=\sum_{k=1}^{3}\alpha_{t}^{(k)}\cdot H_{t}^{(k)},\quad\alpha_{t}^{(k)}=\frac{\exp(w^{\top}\tanh(w_{f}H_{t}^{(k)}))}{\sum_{j}\exp(w^{\top}\tanh(w_{f}H_{t}^{(j)}))}$ (15)
The fused representation $H_{t}$ serves as input for subsequent generation modules.
Confidence-Aware Causal Attention Mechanism
To improve robustness in autoregressive pose generation, we propose a Confidence-Aware Causal Attention module (Figure 3), which integrates keypoint confidence scores into the attention computation to downweight unreliable inputs.
Given a sentence $S$, the model generates a pose sequence $x={x^{(1)},\dots,x^{(T)}}$, where each $x^{(t)}\in\mathbb{R}^{J\times d}$ represents $J$ keypoints. Each keypoint has an associated confidence score $c_{i}^{(t)}\in[0,1]$.
Table 1: Performance comparison on PHOENIX14T
Methods DEV TEST
[…]
PT ptslp 12.51 3.88 11.87 96.85 - - 13.35 4.31 13.17 96.50 - -
G2P-DDM g2pddm - - - - - - 16.11 7.50 - 77.26 - -
GCDM gcdm 22.88 7.64 23.35 82.81 11.18 39.87 22.03 7.91 23.20 81.94 11.10 49.22
GEN-OBT gen-obt 24.92 8.68 25.21 82.36 - - 23.08 8.01 23.49 81.78 - -
Sign-IDD signidd 25.40 8.93 27.60 77.72 5.09 39.11 24.80 9.08 26.58 76.66 6.20 47.19
HybridSign (Ours) 26.98 9.26 28.07 75.81 4.07 38.28 25.77 10.03 27.97 75.02 4.96 45.50
Ground Truth 29.77 12.13 29.60 74.17 0.00 0.00 29.76 11.93 28.98 71.94 0.00 0.00
Table 2: Performance comparison on How2Sign
Methods DEV TEST
[…]
PT ptslp 14.34 4.09 8.12 96.91 10.53 55.02 14.05 4.17 8.42 96.47 10.18 54.57
G2P-DDM g2pddm 19.82 5.37 12.47 90.05 8.25 50.48 19.48 5.12 12.28 89.58 7.97 49.83
GCDM gcdm 26.43 5.84 15.53 91.92 6.32 46.19 25.91 5.57 15.21 91.43 6.13 45.71
GEN-OBT gen-obt 28.63 6.14 16.23 91.07 7.08 48.03 27.82 5.92 15.88 90.63 6.87 47.28
Sign-IDD signidd 29.12 6.27 17.01 89.95 4.76 35.03 28.90 6.06 16.21 89.98 4.86 39.02
HybridSign (Ours) 30.71 6.92 18.96 87.11 3.76 34.19 30.12 6.48 18.02 88.30 3.89 37.10
Ground Truth 35.20 8.89 22.45 83.79 0.00 0.00 34.01 8.03 21.87 81.94 0.00 0.00
Confidence-Weighted Attention.
In causal attention, timestep $t$ only attends to $1,\dots,t$}:
$\alpha_{t,s}=\frac{\exp\left(\frac{Q_{t}K_{s}^{\top}}{\sqrt{d_{k}}}\right)}{\sum_{j=1}^{t}\exp\left(\frac{Q_{t}K_{j}^{\top}}{\sqrt{d_{k}}}\right)}$ (16)
We introduce a confidence bias to this formulation:
$\alpha_{t,s}=\frac{\exp\left(\frac{Q_{t}K_{s}^{\top}}{\sqrt{d_{k}}}+\beta\cdot\bar{c}^{(s)}\right)}{\sum_{j=1}^{t}\exp\left(\frac{Q_{t}K_{j}^{\top}}{\sqrt{d_{k}}}+\beta\cdot\bar{c}^{(j)}\right)}$ (17)
where $\bar{c}^{(s)}=\frac{1}{J}\sum_{i=1}^{J}c_{i}^{(s)}$ is the average keypoint confidence, and $\beta$ is a learnable scalar controlling the strength of the bias.
This allows the model to attend more to reliable frames during decoding.
Training Objectives
Our model is optimized using a composite loss that combines three complementary objectives: joint accuracy, kinematic consistency, and temporal alignment.
Joint loss
To encourage accurate joint localization, we minimize the mean absolute error between predictions and ground truth:
$\mathcal{L}_{\mathrm{joint}}\;=\;\frac{1}{J}\sum_{j=1}^{J}\|p_{j}-\hat{p}_{j}\|_{1},$ (18)
where $p_{j}\in\mathbb{R}^{d}$ and $\hat{p}_{j}\in\mathbb{R}^{d}$ denote the ground-truth and predicted position of joint $j$, and $\|\cdot\|_{1}$ is the $\ell_{1}$ norm.
Bone loss
To preserve bone orientations and kinematic consistency, we penalize orientation deviations:
$\mathcal{L}_{\mathrm{bone}}\;=\;\frac{1}{B}\sum_{b=1}^{B}\|q_{b}-\hat{q}_{b}\|_{2}^{2},$ (19)
where $q_{b}$ and $\hat{q}_{b}$ are the ground-truth and predicted orientation (e.g. unit quaternions or axis-angle vectors) of bone $b$, and $\|\cdot\|_{2}$ denotes the Euclidean norm. If orientations are represented as quaternions, consider using a geodesic/angular metric as an alternative.
Soft-DTW loss
To align predicted and ground-truth pose sequences temporally, we employ differentiable Dynamic Time Warping (Soft-DTW) softdtw :
$\mathcal{L}_{\mathrm{soft\text{-}dtw}}\;=\;\min\nolimits_{A}^{\gamma}\sum_{(i,j)\in A}\|x_{i}-y_{j}\|_{2}^{2},$ (20)
where $x_{i}$ and $y_{j}$ are per-frame pose descriptors, $A$ is an alignment path, and $\gamma>0$ controls the smoothness of the soft minimum:
$\min\nolimits^{\gamma}\{a_{1},\dots,a_{k}\}=-\gamma\log\!\Bigg(\sum_{i=1}^{k}e^{-a_{i}/\gamma}\Bigg).$ (21)
Adaptive total loss
We balance the three terms with dynamic weights computed from the inverse exponential moving average (EMA) of each loss dwa . Let $\bar{\mathcal{L}}_{i}^{(t)}$ be the EMA of loss $i$ at iteration $t$, and $\epsilon>0$ a small constant to avoid division by zero. The weight for loss $i$ is
$\lambda_{i}^{(t)}\;=\;\frac{\big(\bar{\mathcal{L}}_{i}^{(t)}+\epsilon\big)^{-1}}{\sum_{j=1}^{N}\big(\bar{\mathcal{L}}_{j}^{(t)}+\epsilon\big)^{-1}},\qquad i=1,\dots,N,$ (22)
with $N=3$ in our case. The final objective at iteration $t$ is
$\mathcal{L}_{\mathrm{total}}^{(t)}\;=\;\lambda_{1}^{(t)}\,\mathcal{L}_{\mathrm{joint}}^{(t)}\;+\;\lambda_{2}^{(t)}\,\mathcal{L}_{\mathrm{bone}}^{(t)}\;+\;\lambda_{3}^{(t)}\,\mathcal{L}_{\mathrm{soft\text{-}dtw}}^{(t)}.$ (23)
This adaptive weighting scheme emphasizes losses that are relatively smaller (via inverse EMA), promoting stable and balanced training across spatial accuracy, kinematic consistency, and temporal alignment.
Experiment
Figure 4: Visualization examples of generated poses on How2Sign. We compare HybridSign with the ground-truth poses and the original video frames for clear evaluation.
Experimental Setup
Datasets.
We evaluate the proposed method on two benchmarks: PHOENIX14T and How2Sign. The PHOENIX14T phoenix is a German Sign Language (DGS) corpus that consists of 8,257 sentence-level samples, encompassing 2,887 unique German words. The How2Sign how2sign is a large-scale American Sign Language (ASL) corpus designed for sign language understanding and generation tasks. It contains over 80 hours of multimodal data.
Evaluation Metrics.
Following the existing works g2pddm ; gcdm ; signidd ; signd2c , we used a pre-trained SLT model slt for back-translation, converting generated sign language pose sequences back to text, and then evaluated the results with BLEU bleu , ROUGE rouge , WER, DTW, and FID.
Comparisons with Baseline Models
We conducted both qualitative and quantitative analyses to evaluate the model’s performance.
Table 3: Latency and throughput comparison between the baseline models and this model with 3 expert modules for generating 60-frame sign language pose sequence.
[…]
GCDM gcdm 52.18 1.15
Sign-IDD signidd 40.31 1.49
G2P-DDM g2pddm 25.78 2.33
HybridSign 5.90 10.17
Quantitative Analysis.
Tables 1 and 2 present a performance comparison between the proposed HybridSign and various baseline models on the PHOENIX14T phoenix and How2Sign how2sign datasets, respectively. As shown, HybridSign outperforms all baseline models across all evaluation metrics.
Table 3 compares latency and throughput with the baseline models. Here, latency refers to the time required to generate the first frame of the 60-frame sequence, rather than the total time to produce the entire video. This reflects the model’s initial response speed, which is important for real-time applications. HybridSign achieves significantly lower latency and higher throughput than the baselines, indicating its efficiency and practicality in real-time sign language production.
Performance differences across the two datasets are relatively small, likely due to the use of 2D pose sequences. While this may slightly reduce generation quality, it lowers computational complexity and improves training and inference efficiency. To compensate, we incorporate the Multi-Scale Pose Representation Module and the Confidence-Attention Causal Mechanism, which enhance overall performance. Notably, the DTW score is reduced by approximately 20% thanks to the Soft-DTW loss softdtw , which improves temporal alignment.
Qualitative Analysis.
Figure 4 illustrates examples of sign poses generated by HybridSign alongside the corresponding ground truth. While minor differences are observable, the generated poses remain highly similar overall and do not hinder interpretability. This further supports the effectiveness and practicality of our approach.
From the generated results, there are still some subtle differences compared to the ground truth, such as variations in bone lengths or joint bending angles. However, based on the results shown in the table 1 and 2, after performing back translation on the sign pose sequences, the B1 and B4 scores on the two datasets are 26.98, 9.26 and 30.71, 6.92 respectively. Therefore, we believe these slight differences are acceptable.
Table 4: Ablation study results of modes on the How2Sign dataset.
[…]
Diffusion Mode 30.25 6.55 8.06 32.89 1.83
Autoregressive Mode 26.15 5.40 4.49 5.53 10.85
Hybrid Mode 30.12 6.48 3.89 5.90 10.17
Table 5: Ablation study results of modules on the How2Sign dataset. CA denotes Causal Attention, CACA denotes Confidence-Aware Causal Attention.
[…]
RNN 23.47 5.02 6.89 4.72 9.71
CA 25.08 5.83 5.50 5.48 10.95
CACA 30.12 6.48 3.89 5.90 10.17
Table 6: Ablation study results of expert numbers on the How2Sign dataset.
[…]
1 (whole pose) 22.33 5.14 7.02 7.69 7.80
4 (face + body + lh + rh) 29.17 6.03 5.72 7.73 7.76
3 (face + body + hands) 30.12 6.48 3.89 5.90 10.17
Ablation Study
We conducted three ablation studies on the generation of 60-frame sign language pose sequences to systematically evaluate our method. Specifically, we compared the proposed hybrid approach with a pure diffusion model and a pure autoregressive model, and further investigated the impact of different attention mechanisms and varying numbers of expert modules.
Table  4 compares the performance and efficiency of a pure diffusion model and a pure autoregressive model. Specifically, the diffusion model adopts the DiT architecture  dit , while the autoregressive model uses a Transformer decoder with teacher forcing  attention . Experimental results show that, thanks to its multi-step denoising process, the diffusion model achieves very high generation quality and reaches the best scores on both the B1 and B4 metrics. However, its temporal consistency and generation efficiency lag significantly behind the other models, making it unsuitable for real-time scenarios. In contrast, the autoregressive model generates frame-by-frame, resulting in extremely low first-frame latency, but suffers from a noticeable quality drop due to the mismatch between training and inference. Our proposed approach combines the high-quality generation of diffusion models with the frame-by-frame generation of autoregressive models, achieving both strong performance and efficiency.
Table 5 presents the performance differences among various autoregressive implementations, including a recurrent neural network (RNN) rnn , a standard causal attention mechanism attention , and a confidence-aware causal attention mechanism. The results indicate that the RNN-based model achieves very low generation quality, but due to its computational simplicity compared to attention mechanisms, it exhibits excellent first-frame latency. However, its strict sequential dependency limits overall efficiency, resulting in lower throughput than attention-based methods. In contrast, the confidence-aware causal attention mechanism incorporates confidence scores, assigning different attention weights based on the reliability of keypoints, which leads to the best generation quality while maintaining strong applicability in real-time scenarios.
Table 6 presents the results of varying the number of expert modules. The experiments include using a single expert module (treating the entire pose as a whole), using three expert modules to handle the face, body, and hands separately, and using four expert modules to handle the face, body, left hand, and right hand individually. The results show that with a single expert, the model cannot capture local details, leading to a significant drop in generation quality. Moreover, since the single module cannot take advantage of parallel computation, its efficiency is relatively low. Using four experts increases inter-module dependencies and the difficulty of feature fusion, which can cause information conflicts or redundancy and negatively affect generation quality. Although parallelization eliminates most of the serial computation overhead, the integration of outputs from multiple modules still requires some serial processing, limiting overall performance. In contrast, using three expert modules strikes the optimal balance, achieving the best trade-off between generation quality and efficiency.
Conclusions
This paper explores a hybrid framework that combines autoregressive and diffusion models for sign language production, enabling real-time generation of individual sign language pose frames. The framework leverages the inherent sequential modeling capability and frame-by-frame output of autoregressive models, along with the high-quality generation strengths of diffusion models. Additionally, it incorporates multi-scale pose representation modeling and a confidence-aware causal attention mechanism, which further improve the naturalness and expressive accuracy of the generated sign language. Extensive experiments on several benchmark datasets demonstrate the effectiveness of the proposed method, particularly in real-time applications.